OpenStudy (anonymous):

Evaluate the expression and write the result in the form a + bi. 8i(1/8 − i)

OpenStudy (anonymous):

\[8i (1/8 - i) = i - 8i^2 = i - 8(-1) = i +8\]

OpenStudy (anonymous):

or 8 + 1i

OpenStudy (anonymous):

i wonder what form my prof expects it in

OpenStudy (anonymous):

a+ bi form would be 8 + i

OpenStudy (anonymous):

put the one in from just in case I guess

OpenStudy (anonymous):

so just put: 8 + 1i